# The finite groups with three automorphism orbits

### v1

@@ 2.2. Special $p$-groups of exponent $p$
Let $N$ be a non-abelian $3$-orbit $p$-group with $p$ an odd prime. Then non-identity elements in the center ${\mathbf{Z}}(N)$ form an automorphism orbit, and elements in $N\setminus{\mathbf{Z}}(N)$ form a different automorphism orbit. It follows that both ${\mathbf{Z}}(N)$ and $N/{\mathbf{Z}}(N)$ are elementary abelian groups, and ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$. So $N$ is a special $p$-group of exponent $p$ or $p^{2}$. If the exponent of $N$ is $p^{2}$, then all elements of $N$ of order $p$ form an automorphism orbit, and so $N$ is abelian by [27, Corollary 3]. This leads to the following lemma.
@@
For a non-abelian special $p$-group $N$, ${\mathbf{Z}}(N)=N^{\prime}=\Phi(N)$ and $N/\Phi(N)$ are elementary abelian $p$-groups, and thus they can be viewed as vector spaces over field ${\mathbb{F}}_{p}$. Then the actions of $A:={\mathrm{Aut}}(N)$ on $\Phi(N)$ and $N/\Phi(N)$ are linear actions. It yields that $\Phi(N)$ and $N/\Phi(N)$ can be viewed as ${\mathbb{F}}_{p}A$-modules. For elements $x,y,z\in N$, we have $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as ${\mathbf{Z}}(N)=N^{\prime}$. Thus, the commutator operator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] on $N$ resembles an alternating bilinear form. Let $V$ be an ${\mathrm{F}}G$-module. The exterior square $\Lambda_{\mathrm{F}}^{2}(V)$ is the quotient ${\mathrm{F}}G$-module $(V\otimes V)/S$, where $S$ is spanned by vectors $v\otimes v$ for all $v\in V$. The exterior square has a universal property associated with alternating bilinear form. The following lemma is given in [22, Lemma 3.1].
@@
Let $N$ be a non-abelian special $p$-group with $p$ prime such that $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{n}$. Let $K$ and $L$ be the kernels of $A:={\mathrm{Aut}}(N)$ acting on $V:=N/{\mathbf{Z}}(N)$ and $M:={\mathbf{Z}}(N)$, respectively. Let $G=A/K$. Then the following statements hold.
@@
    $N$ is an $r$-group with $r=o(V)+o(M)-1$, where $o(V)$ and $o(M)$ are numbers of $G$-orbits on $V$ and $M$, respectively.
@@
Recall that, for an odd prime $p$, a non-abelian $3$-orbit $p$-group $N$ is a special $p$-group of exponent $p$. Let ${\mathcal{H}}_{n,p}$ denote the $n$-generator free group in the variety of groups that have exponent $p$ and nilpotency class $2$. Then $N$ is a quotient of ${\mathcal{H}}_{n,p}$ when $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. We remark that ${\mathcal{H}}_{n,p}$ is the so-called the universal generalized Heisenberg group in [24], and it is a finite special $p$-group with $|{\mathbf{Z}}({\mathcal{H}}_{n,p})|=p^{\binom{n}{2}}$ (see [3, Page 107]). We formulate an equivalent definition below.
@@
${\mathcal{H}}_{n,p}:=\langle\alpha_{1},...,\alpha_{n}|\alpha_{i}^{p}=[\alpha_{i},[\alpha_{j},\alpha_{k}]]=1\mbox{ for $i,j,k=1,...,n$}.\rangle$
Let ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and let $W<{\mathbf{Z}}({\mathcal{H}})$. The automorphism group of ${\mathcal{H}}/W$ for $W<{\mathbf{Z}}({\mathcal{H}})$ is described in [13, Lemma 3] which is based on [26, Theorem 2.10]. We state the results below.
@@
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Identify $V:=N/{\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, and identify ${\mathbf{Z}}(N)$ with ${\mathbf{Z}}({\mathcal{H}})/W$. Then the following statements hold.
@@
    $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{Aut}}(V)={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$, and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as an ${\mathbb{F}}_{p}X$-module.
@@
    ${\mathrm{Aut}}(N)^{V}\leqslant X$ is equal to $X_{W}$, the maximal subgroup of $X$ stabilizes the subspace $W$ of ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
Using the notations above, Lemma 2.8 (i) deduces that
@@
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}:={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V:=N/{\mathbf{Z}}(N)$, and let $K$ be the kernel of ${\mathrm{Aut}}(N)$ acting on $V$. Then the group extension ${\mathrm{Aut}}(N)=K.{\mathrm{Aut}}(N)^{V}$ is split.
@@
First, we consider $N={\mathcal{H}}$. Then $K\cong p^{\frac{n^{2}(n-1)}{2}}$ and ${\mathrm{Aut}}({\mathcal{H}})^{V}\cong{\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ by Lemma 2.10. Recall that ${\mathcal{H}}=\langle\alpha_{1},...,\alpha_{n}\rangle$. Let $k\in{\mathbb{F}}_{p}^{\times}$ be a primitive element, then there exists $\psi\in{\mathrm{Aut}}({\mathcal{H}})$ of order $p-1$ such that
@@
Let ${\mathcal{H}}:={\mathcal{H}}_{n,p}$, and let $W_{1},W_{2}$ be two subgroups of ${\mathbf{Z}}({\mathcal{H}})$. View $W_{1}$ and $W_{2}$ as two subspaces of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ where $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then ${\mathcal{H}}/W_{1}\cong{\mathcal{H}}/W_{2}$ if there exists $g\in{\mathrm{GL}}(V)$ such that $W_{1}^{g}=W_{2}$ in $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
@@ 2.3. Extraspecial $q$-groups
Let $p$ be an odd prime. We introduce a new family of special $p$-groups for odd prime $p$ in this part. Recall that the extraspecial $p$-group of exponent $p$, denoted by $p^{1+2m}_{+}$, is the central product of $m$ copies of ${\mathrm{SL}}(3,p)_{p}$ for odd prime $p$. This leads us to start with the Sylow $p$-subgroups of ${\mathrm{SL}}(3,p^{n})$.
@@
Let $N:=q^{1+2}_{+}$ for $q=p^{n}$ where $p$ is an odd prime. Then $p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)\lesssim{\mathrm{Aut}}(N)$ and $N$ is a $3$-orbit group.
@@
$A(x):=\begin{pmatrix}1&x&0\\ 0&1&0\\ 0&0&1\end{pmatrix},B(y):=\begin{pmatrix}1&0&0\\ 0&1&y\\ 0&0&1\end{pmatrix}\mbox{ and }C(z):=\begin{pmatrix}1&0&z\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
Since $N$ is a special $p$-group of exponent $p$, it is isomorphic to ${\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{2n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $\alpha(x)$ and $\beta(y)$ be preimages of $A(x)$ and $B(y)$ in ${\mathcal{H}}$ for $x,y\in{\mathbb{F}}_{q}$, respectively. Hence, we have
${\mathcal{H}}=\langle\alpha(x),\beta(y):x,y\in{\mathbb{F}}_{q}\rangle.$
By Lemma 2.10 (i), $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2n,p)$ where $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$. Note that $V$ can be viewed as a $2$-dimensional ${\mathbb{F}}_{q}$-space $V_{q}$ with basis $\alpha(1){\mathbf{Z}}({\mathcal{H}})$ and $\beta(1){\mathbf{Z}}({\mathcal{H}})$. Let $G\cong{\Gamma\mathrm{L}}(2,p^{n})$ be the subgroup of $X$ acting naturally on $V_{q}$. Now we prove that $W^{g}=W$ for $g\in G$, and then $G\lesssim{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Note that each element in ${\mathbf{Z}}({\mathcal{H}})={\mathcal{H}}^{\prime}$ is a product of elements of forms $[\alpha(x),\alpha(y)]$, $[\beta(x),\beta(y)]$ and $[\alpha(x),\beta(y)]$ for $x,y\in{\mathbb{F}}_{q}$.
  1. (1)
@@
    $\prod_{i=1}^{k}[\alpha(x_{i}),\beta(y_{i})]^{g}=\prod_{i=1}^{k}[\alpha(x_{i}a)\beta(x_{i}b),\alpha(y_{i}c)\beta(y_{i}d)]=\prod_{i=1}^{k}s_{i}\cdot\prod_{i=1}^{k}t_{i}.$
    where $s_{i}=[\alpha(x_{i}a),\alpha(y_{i}c)][\beta(x_{i}b),\beta(y_{i}d)]$, and $t_{i}=[\alpha(x_{i}a),\beta(y_{i}d)][\beta(x_{i}b),\alpha(y_{i}c)]$. By (1) and (2), $s_{i}$’s are in $W$. The image of $\prod_{i=1}^{k}t_{i}$ in $N$ is equal to
@@
Note that ${\Gamma\mathrm{L}}(2,q)$ contains the Singer cyclic which is a cyclic group of order $q^{2}-1$ acting regularly on non-identity elements of ${\mathbb{F}}_{q}^{2}$. This yields a solvable subgroup of ${\mathrm{Aut}}(N)$ which acts on $N$ with $3$ orbits. Immediately, we obtain the follow lemma, which is Theorem C (ii).
@@
Let $N:=q^{1+2}_{+}$, and let $A\cong p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)$ be the subgroup of ${\mathrm{Aut}}(N)$ given in Lemma 2.14. Then there exists $\varphi\in A$ of order $q^{2}-1$ such that $\varphi$ induces a transitive action on non-identity elements of $N/{\mathbf{Z}}(N)$. Let $\xi\in{\mathrm{Aut}}(N)$ induced by conjugating the diagonal matrix $\mathrm{diag}(\mu,1,1)$ for some generator $\mu$ of ${\mathbb{F}}_{q}^{\times}$. We remark that $\xi\in A$ and $\xi$ acts transitively on non-identity elements of ${\mathbf{Z}}(N)$. Let $K\cong p^{2n^{2}}$ be the kernel of $A$ acting on $N/{\mathbf{Z}}(N)$. Then $\langle K,\varphi,\xi\rangle$ is a solvable subgroup of $A\leqslant{\mathrm{Aut}}(N)$, and it acts on elements of $N$ with exactly $3$ orbits by Lemma 2.8 (iii). Thus, ${\mathrm{Aut}}(N)$ has a solvable subgroup with exactly $3$ orbits on $N$.
Note that $H:=\langle M(a,b)\mid a,b\in{\mathbb{F}}_{q^{2}},b+b^{q}+a^{1+q}=0\rangle$ is a Sylow $p$-subgroup of ${\mathrm{SU}}(3,q)$ (see [11, page 249]), where
$M(a,b)=\begin{pmatrix}1&a&b\\ 0&1&a^{q}\\ 0&0&1\end{pmatrix}.$
@@
Hence, $D$ induces an automorphism $\xi^{\prime}\in{\mathrm{Aut}}(H)$ such that $\langle\xi^{\prime}\rangle$ acting transitively on non-identity elements of both $H/{\mathbf{Z}}(H)$ and ${\mathbf{Z}}(H)$. By Lemma 2.8 (i) and (iii), ${\mathrm{Aut}}(H)$ has a solvable subgroup with exactly $3$ orbits on $H$.
@@
$q^{1+2m}_{+}:=q^{1+2}_{+}\circ\cdots\circ q^{1+2}_{+}=(q^{1+2}_{+})^{m}/C,$
@@
It is straightforward to see that $q^{1+2m}_{+}$ is a non-abelian special $p$-group of exponent $p$ such that ${\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{n}$ and $q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{2mn}$. Let $\pi_{i}$ be the nature homomorphism from $q^{1+2}_{+}$ to the image of the $i$-th component in $q^{1+2m}_{+}$ for $i=1,...,m$. Remark that each matrix $Z\in{\mathbf{Z}}(q^{1+2}_{+})$ satisfies $\pi_{i}(Z)=\pi_{j}(Z)$ for $i,j=1,...,m$. Then we define the following notations for $i=1,...,m$ and $a,b,c\in{\mathbb{F}}_{q}$:
$E_{i}(a):=\pi_{i}\begin{pmatrix}1&a&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\mbox{, }F_{i}(b):=\pi_{i}\begin{pmatrix}1&0&0\\ 0&1&b\\ 0&0&1\end{pmatrix}\mbox{, and }Z(c)=\pi_{i}\begin{pmatrix}1&0&c\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
Let $N:=q^{1+2m}_{+}$ with $q=p^{n}$ for odd prime $p$. Then ${\mathrm{Aut}}(N)$ has a subgroup $A$ such that $A^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and $A^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$, where $V=N/{\mathbf{Z}}(N)$.
@@
Since $N$ is a non-abelian special $p$-group of exponent $p$, $N={\mathcal{H}}/W$ with ${\mathcal{H}}:={\mathcal{H}}_{2mn,p}$ and $W\leqslant\Phi({\mathcal{H}})$. Identify $V:=N/{\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, and let $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ as in Lemma 2.10. For convenience, let $\overline{\alpha}:=\alpha{\mathbf{Z}}({\mathcal{H}})\in V$ for $\alpha\in{\mathcal{H}}$. Let $e_{i}(a)$ and $f_{j}(b)$ be preimages of $E_{i}(a)$ and $F_{i}(b)$ in ${\mathcal{H}}$, respectively. Hence, $V$ can be viewed as an ${\mathbb{F}}_{q}$-symplectic space $V_{q}$ equipped with the symplectic form ${\mathcal{B}}$ and a basis
@@
Step 1: we show that the image of $[\alpha,\beta]$ in $N$ is $Z\left({\mathcal{B}}\left(\overline{\alpha},\overline{\beta}\right)\right)$ for any $\alpha,\beta\in{\mathcal{H}}$. Assume that
@@
where $a_{i},b_{i},c_{i},d_{i}\in{\mathbb{F}}_{q}$. Then the image of $[\alpha,\beta]$ in $N$ equals to

### v2

@@ 2.2. Special $p$-groups of exponent $p$
Let $N$ be a non-abelian $3$-orbit $p$-group with $p$ an odd prime. Then non-identity elements in the center ${\mathbf{Z}}(N)$ form an automorphism orbit, and elements in $N\setminus{\mathbf{Z}}(N)$ form a different automorphism orbit. It follows that both ${\mathbf{Z}}(N)$ and $N/{\mathbf{Z}}(N)$ are elementary abelian groups, and ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$. So $N$ is a special $p$-group of exponent either $p$ or $p^{2}$. If the exponent of $N$ is $p^{2}$, then all elements of $N$ of order $p$ form an automorphism orbit, and so $N$ is abelian by [32, Corollary 3]. This leads to the following lemma.
@@
For a non-abelian special $p$-group $N$, ${\mathbf{Z}}(N)=N^{\prime}=\Phi(N)$ and $N/\Phi(N)$ are elementary abelian $p$-groups, and thus they can be viewed as vector spaces over field ${\mathbb{F}}_{p}$. Let $A={\mathrm{Aut}}(N)$. Then the actions of $A$ on $\Phi(N)$ and $N/\Phi(N)$ are linear actions. It yields that $\Phi(N)$ and $N/\Phi(N)$ can be viewed as ${\mathbb{F}}_{p}A$-modules. For elements $x,y,z\in N$, we have $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as ${\mathbf{Z}}(N)=N^{\prime}$. Thus, the commutator operator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] on $N$ resembles an alternating bilinear form. Let $V$ be an ${\mathrm{F}}G$-module. The exterior square $\Lambda_{\mathrm{F}}^{2}(V)$ is the quotient ${\mathrm{F}}G$-module $(V\otimes V)/S$, where $S$ is spanned by vectors $v\otimes v$ for all $v\in V$. The exterior square has a universal property associated with alternating bilinear form. The following lemma is given in [25, Lemma 3.1].
@@
Let $N$ be a non-abelian special $p$-group with $p$ prime such that $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{n}$. Set $V=N/{\mathbf{Z}}(N)$ and $M={\mathbf{Z}}(N)$. Let $K$ and $L$ be the kernels of ${\mathrm{Aut}}(N)$ acting on $V$ and $M$, respectively. Define $G=A/K$. Then the following statements hold.
@@
    $N$ is an $r$-orbit group with $r=o(V)+o(M)-1$, where $o(V)$ and $o(M)$ are numbers of $G$-orbits on $V$ and $M$, respectively.
@@
Recall that, for an odd prime $p$, a non-abelian $3$-orbit $p$-group $N$ is a special $p$-group of exponent $p$. Let ${\mathcal{H}}_{n,p}$ denote the $n$-generator free group in the variety of groups that have exponent $p$ and nilpotency class $2$. Then $N$ is a quotient of ${\mathcal{H}}_{n,p}$ when $N/\Phi(N)\cong{\mathbb{Z}}_{p}^{n}$. We remark that ${\mathcal{H}}_{n,p}$ is the so-called the universal generalized Heisenberg group in [27], and it is a finite special $p$-group with $|{\mathbf{Z}}({\mathcal{H}}_{n,p})|=p^{\binom{n}{2}}$ (see [4, Page 107]). We formulate an equivalent definition below.
@@
${\mathcal{H}}_{n,p}=\langle\alpha_{1},...,\alpha_{n}|\alpha_{i}^{p}=[\alpha_{i},[\alpha_{j},\alpha_{k}]]=1\mbox{ for $i,j,k=1,...,n$}.\rangle$
Let ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and let $W<{\mathbf{Z}}({\mathcal{H}})$. The automorphism group of ${\mathcal{H}}/W$ for $W<{\mathbf{Z}}({\mathcal{H}})$ is described in [15, Lemma 3] which is based on [29, Theorem 2.10]. We state the results as below.
@@
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$, and let $M={\mathbf{Z}}(N)={\mathbf{Z}}({\mathcal{H}})/W$. Define $X={\mathrm{Aut}}({\mathcal{H}})^{V}$ and $G={\mathrm{Aut}}(N)^{V}$. Then the following statements hold.
@@
    $X={\mathrm{Aut}}(V)={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$, and ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as an ${\mathbb{F}}_{p}X$-module.
@@
    $G\leqslant X$ is equal to $X_{W}$, the maximal subgroup of $X$ that stabilizes the subspace $W$ of ${\mathbf{Z}}({\mathcal{H}})$.
Using the notations above, Lemma 2.8 (i) implies that
@@
Suppose that $N={\mathcal{H}}/W$ where ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V=N/{\mathbf{Z}}(N)$, and let $K$ be the kernel of ${\mathrm{Aut}}(N)$ acting on $V$. Then the group extension ${\mathrm{Aut}}(N)=K.{\mathrm{Aut}}(N)^{V}$ is split.
@@
First, we consider $N={\mathcal{H}}$. Then $K\cong p^{\frac{n^{2}(n-1)}{2}}$ and ${\mathrm{Aut}}({\mathcal{H}})^{V}\cong{\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$ by Lemma 2.10. Recall that ${\mathcal{H}}=\langle\alpha_{1},...,\alpha_{n}\rangle$. Let $k\in{\mathbb{F}}_{p}^{\times}$ be a primitive element. Then there exists $\psi\in{\mathrm{Aut}}({\mathcal{H}})$ of order $p-1$ such that
@@
Let ${\mathcal{H}}={\mathcal{H}}_{n,p}$, and let $W_{1},W_{2}$ be two subgroups of ${\mathbf{Z}}({\mathcal{H}})$. View $W_{1}$ and $W_{2}$ as two subspaces of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ where $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then ${\mathcal{H}}/W_{1}\cong{\mathcal{H}}/W_{2}$ if there exists $g\in{\mathrm{GL}}(V)$ such that $W_{1}^{g}=W_{2}$ in $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
@@ 2.3. Extraspecial $q$-groups
Let $p$ be an odd prime. We introduce a new family of special $p$-groups in this part. Recall that the extraspecial $p$-group of exponent $p$, denoted by $p^{1+2m}_{+}$, is the central product of $m$ copies of ${\mathrm{SL}}(3,p)_{p}$. This leads us to start with the Sylow $p$-subgroups of ${\mathrm{SL}}(3,p^{n})$.
@@
Let $N=q^{1+2}_{+}$ for $q=p^{n}$ where $p$ is an odd prime. Then $p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)\lesssim{\mathrm{Aut}}(N)$ and $N$ is a $3$-orbit group.
@@
$A(x)=\begin{pmatrix}1&x&0\\ 0&1&0\\ 0&0&1\end{pmatrix},B(y)=\begin{pmatrix}1&0&0\\ 0&1&y\\ 0&0&1\end{pmatrix}\mbox{ and }C(z)=\begin{pmatrix}1&0&z\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
Since $N$ is a special $p$-group of exponent $p$, it is isomorphic to ${\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{2n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $\alpha(x)$ and $\beta(y)$ be preimages of $A(x)$ and $B(y)$ in ${\mathcal{H}}$ for $x,y\in{\mathbb{F}}_{q}$, respectively. Note that each element in $N$ can be expressed as a product of elements $A(x)$, $B(y)$ and $C(z)$ for some $x,y,z\in{\mathbb{F}}_{q}$. With the equalities in (a), we have that $C(z)\in\langle A(x),B(y):x,y\in{\mathbb{F}}_{q}\rangle$. This yields that
$N=\langle A(x),B(y):x,y\in{\mathbb{F}}_{q}\rangle\mbox{ and }{\mathcal{H}}=\langle\alpha(x),\beta(y):x,y\in{\mathbb{F}}_{q}\rangle.$
Let $X={\mathrm{Aut}}({\mathcal{H}})^{V}$ with $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. By Lemma 2.10 (i), $X={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2n,p)$. Note that $V$ can be viewed as a $2$-dimensional ${\mathbb{F}}_{q}$-space $V_{q}$ with basis $\alpha(1){\mathbf{Z}}({\mathcal{H}})$ and $\beta(1){\mathbf{Z}}({\mathcal{H}})$. Let $G\cong{\Gamma\mathrm{L}}(2,p^{n})$ be the subgroup of $X$ acting naturally on $V_{q}$.
Now we prove that $W^{g}=W$ for $g\in G$, and then $G\lesssim{\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Note that each element in ${\mathbf{Z}}({\mathcal{H}})={\mathcal{H}}^{\prime}$ is a product of elements of forms $[\alpha(x),\alpha(y)]$, $[\beta(x),\beta(y)]$ and $[\alpha(x),\beta(y)]$ for $x,y\in{\mathbb{F}}_{q}$.
  1. (i)
@@
    $\prod_{i=1}^{k}[\alpha(x_{i}),\beta(y_{i})]^{g}=\prod_{i=1}^{k}[\alpha(x_{i}a)\beta(x_{i}b),\alpha(y_{i}c)\beta(y_{i}d)]=\prod_{i=1}^{k}s_{i}\cdot\prod_{i=1}^{k}t_{i},$
    where $s_{i}=[\alpha(x_{i}a),\alpha(y_{i}c)][\beta(x_{i}b),\beta(y_{i}d)]$, and $t_{i}=[\alpha(x_{i}a),\beta(y_{i}d)][\beta(x_{i}b),\alpha(y_{i}c)]$. By (i) and (ii), each $s_{i}$ lies in $W$. The image of $\prod_{i=1}^{k}t_{i}$ in $N$ is equal to
@@
Note that ${\Gamma\mathrm{L}}(2,q)$ contains the Singer cyclic which is a cyclic group of order $q^{2}-1$ acting regularly on non-identity elements of ${\mathbb{F}}_{q}^{2}$. This yields a solvable subgroup of ${\mathrm{Aut}}(N)$ which acts on $N$ with $3$ orbits. Immediately, we obtain the following lemma, which is Theorem C (ii).
@@
Let $N=q^{1+2}_{+}$, and let $A\cong p^{2n^{2}}{:}{\Gamma\mathrm{L}}(2,q)$ be the subgroup of ${\mathrm{Aut}}(N)$ given in Lemma 2.14. Then there exists $\varphi\in A$ of order $q^{2}-1$ such that $\varphi$ induces a transitive action on non-identity elements of $N/{\mathbf{Z}}(N)$. Let $\xi\in{\mathrm{Aut}}(N)$ induced by conjugating the diagonal matrix $\mathrm{diag}(\mu,1,1)$ for some generator $\mu$ of ${\mathbb{F}}_{q}^{\times}$. We remark that $\xi\in A$ and $\xi$ acts transitively on non-identity elements of ${\mathbf{Z}}(N)$. Let $K\cong p^{2n^{2}}$ be the kernel of $A$ acting on $N/{\mathbf{Z}}(N)$. Then $\langle K,\varphi,\xi\rangle$ is a solvable subgroup of $A\leqslant{\mathrm{Aut}}(N)$, and it acts on elements of $N$ with exactly $3$ orbits by Lemma 2.8 (iii). Thus, ${\mathrm{Aut}}(N)$ has a solvable subgroup with exactly $3$ orbits on $N$.
Let $H$ be a Sylow $p$-subgroup of ${\mathrm{SU}}(3,q)$ generated by matrices $M(a,b)\in{\mathrm{GL}}(3,q^{2})$ such that $b+b^{q}+a^{1+q}=0$ and
[…]
@@
Hence, $D$ induces an automorphism $\xi^{\prime}\in{\mathrm{Aut}}(H)$ such that $\langle\xi^{\prime}\rangle$ acts transitively on non-identity elements of both $H/{\mathbf{Z}}(H)$ and ${\mathbf{Z}}(H)$. By Lemma 2.8 (i) and (iii), ${\mathrm{Aut}}(H)$ has a solvable subgroup with exactly $3$ orbits on $H$.
@@
$q^{1+2m}_{+}=q^{1+2}_{+}\circ\cdots\circ q^{1+2}_{+}=(q^{1+2}_{+})^{m}/C,$
@@
It is straightforward to see that $q^{1+2m}_{+}$ is a non-abelian special $p$-group of exponent $p$ such that ${\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{n}$ and $q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})\cong{\mathbb{Z}}_{p}^{2mn}$. Let $\pi_{i}$ be the natural homomorphism from $q^{1+2}_{+}$ to the image of the $i$-th component in $q^{1+2m}_{+}$ for $i=1,...,m$. Remark that each matrix $Z\in{\mathbf{Z}}(q^{1+2}_{+})$ satisfies $\pi_{i}(Z)=\pi_{j}(Z)$ for $i,j=1,...,m$. Then we define the following notations for $i=1,...,m$ and $a,b,c\in{\mathbb{F}}_{q}$:
$E_{i}(a)=\pi_{i}\begin{pmatrix}1&a&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\mbox{, }F_{i}(b)=\pi_{i}\begin{pmatrix}1&0&0\\ 0&1&b\\ 0&0&1\end{pmatrix}\mbox{, and }Z(c)=\pi_{i}\begin{pmatrix}1&0&c\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
Let $N=q^{1+2m}_{+}$ with $q=p^{n}$ for odd prime $p$. Then ${\mathrm{Aut}}(N)$ has a subgroup $A$ such that $A^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and $A^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$, where $V=N/{\mathbf{Z}}(N)$.
@@
Since $N$ is a non-abelian special $p$-group of exponent $p$, $N={\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{2mn,p}$ and $W\leqslant\Phi({\mathcal{H}})$. Identify $V=N/{\mathbf{Z}}(N)$ with ${\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$, and let $X={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)$ as in Lemma 2.10. For convenience, let $\overline{\alpha}=\alpha{\mathbf{Z}}({\mathcal{H}})\in V$ for $\alpha\in{\mathcal{H}}$. Let $e_{i}(a)$ and $f_{j}(b)$ be preimages of $E_{i}(a)$ and $F_{i}(b)$ in ${\mathcal{H}}$, respectively. Hence, $V$ can be viewed as an ${\mathbb{F}}_{q}$-symplectic space $V_{q}$ equipped with the symplectic form ${\mathcal{B}}$ and a basis
@@
Step 1: we show that the image of $[\alpha,\beta]$ in $N$ is equal to $Z\left({\mathcal{B}}\left(\overline{\alpha},\overline{\beta}\right)\right)$ for any $\alpha,\beta\in{\mathcal{H}}$. Assume that
@@
where $a_{i},b_{i},c_{i},d_{i}\in{\mathbb{F}}_{q}$. Then the image of $[\alpha,\beta]$ in $N$ equals